# Design of a CsI(Tl) Calorimeter for Muonium-to-Antimuonium Conversion Experiment

### v3

@@ I Introduction
In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18, 19]. With the construction of the China advanced NUclear physics research Facility (CNUF) and other local muon sources [20, 21, 22, 23], MACE is expected to be the first muon-CLFV experiment in China, offering unique perspectives compared to other related experiments, including those at colliders. The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Positron Transport System (PTS), a Microchannel Plate (MCP) [24, 25], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence. The specifications of muon beamline in MACE is expected to produce a surface muon flux on target of $10^{8}~{}\mu^{+}/$s, with a repetition rate of 30–50 kHz and a pulse width duty cycle below 20%. A schematic of the MACE apparatus is shown in Fig. 1.
@@
In the MACE concept, surface muons with a momentum of 26.3 MeV$/c$ are stopped at a production target where M is formed and then possibly diffuses into the vacuum and converts into $\bar{\text{M}}$. The decay products of $\bar{\text{M}}$ include a Michel electron with an energy up to 52.8 MeV and a positron with atomic-shell energy ($\langle E_{\text{k}}\rangle=13.5$ eV). The Michel electron will be reconstructed by MMS which is expected to achieve $10^{-4}$ level suppression of background in the ECAL end through preliminary simulation. On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the PTS. The transverse position of the $e^{+}$ is conserved during the transportation to MCP, with a flight time FWHM of 6.9 ns. The cylindrical magnet with a field strength of 0.1 T outside ECAL is used to constrain the positron beam spot after transportation. While the MCP, with a timing resolution of 500 ps, is hit by a positron, the ECAL will be triggered and measure the produced two back-to-back annihilation $\gamma$-rays. As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible. Thus, an ECAL with good energy resolution and high efficiency is essential for the design of MACE.
@@ II.1 Overview
Calorimeters with inorganic scintillation crystals generally provide excellent energy resolution and thus have been widely used in particle physics experiments for decades. Commonly used scintillators in calorimeters include NaI(Tl), pure CsI, CsI(Tl), BGO, PbWO4, and others [43]. Table 1 summarizes the properties of these inorganic scintillators. The selection of scintillation crystal for MACE ECAL is based on the energy region of interest less than 1 MeV. NaI(Tl) was widely used in the past decades due to its low cost and high scintillation light output. Nonetheless, it is hygroscopic and has a relatively low density, making it less suitable for large-scale particle physics experiments at present. The problem of hygroscopy is identical for LaBr3(Ce). Moreover, considering the overall budget of this experiment, ultrafast and bright crystals such as LYSO(Ce), GAGG(Ce), or LaBr3(Ce) are not currently included but may be considered for potential future upgrades. The low light output of PbWO4 leads to insufficient energy resolution in the low-energy regime. The same is true for pure CsI. In the MACS experiment, the energy resolution of the pure CsI calorimeter for single $\gamma$ event and double $\gamma$ events was 69% and 52%, respectively [44, 14]. Doping thallium (Tl) in pure CsI can significantly increase its light yield, making it one of the brightest scintillators known. The objective of MACE ECAL is is to achieve excellent energy resolution for $\gamma$-rays produced in $e^{+}e^{-}$ annihilation. This requires optimizing the light yield of the scintillator, which is a key factor in precise energy measurements. Consequently, CsI(Tl) is chosen as the reference scintillator in this ECAL design.
@@ II.2 Optimization of Geometry Parameters
In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the region of positron detection system. Some modules must be removed to ensure that the incoming signal positrons are not obstructed. Specifically, the beam windows are set with an upstream radius of 50 mm and a downstream radius of 5 mm. Consequently, 19 modules at the beam entrance and 1 module at the exit are removed, resulting in a total of 622 modules and achieving 97% coverage of the $4\pi$ solid angle. For an estimation of energy resolution, the class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu. Instead of the fully-covered SD in pervious simulation, the crystal is coupled with an 1.5-inch diameter PMT volume. An isotropic point source, wrapped in 1 mm thick plastic, is placed at the entrance of the module. The energy dependence of FWHM is fit using Eq. 2:
@@
where the best-fit value of $a$, $b$ are $-7.4707$ and $2.7638$, respectively. The fit FWHM and energy resolution results of each full-energy peak are plotted in Fig. 4. For the Type-HEX06 modules, the energy resolution is 10.8% at 511 keV, which is 6–7 times better than the resolution achieved by the calorimeter used in the MACS experiment.
@@ II.3.1 M-to-$\bar{\text{M}}$ conversion signal
Figure 5: Spectra of single Type-PEN module (blue line), single Type-HEX module (green line) and reconstructed annihilation signals (red line) of M-to-$\bar{\text{M}}$ conversion by simulation.
@@ II.3.2 Beam-related backgrounds
The muon beam, produced by stopping $\pi^{+}$ on the surface of a target, often includes positrons or other particles with similar momentum [54]. These background particles can enter the ECAL chamber in the same manner as the signal positrons, potentially causing accidental coincidences between MCP and ECAL. Therefore, it is crucial to estimate their potential contribution.
@@ II.3.3 Cosmic-Ray Muon backgrounds
Cosmic-ray muons are another source of background in ECAL. The EcoMug generator [56] is integrated into the MACE offline software to simulate cosmic-ray muon events on various types of surfaces (plane, cylinder, and half-sphere) with accurate angle and momentum distributions. To evaluate cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4. A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and 10 m length. ECAL is positioned at the center of the tunnel, surrounded by a copper solenoid and an iron shield. Cosmic-ray muons are emitted from a $50\times 50~{}\text{m}^{2}$ plane source at sea level. With a rate of 129 Hz, cosmic-ray muon events equivalent to 1 hour of exposure are generated.
The simulation results indicate that cosmic-ray muons still contribute to the background in the ECAL even after passing through the concrete and shielding. This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays. After applying coincidence selection, no events remain, resulting in background rates of less than 2/2/2 /yr when considering coincidences with the Magnetic Spectrometer. Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 57].
@@ III.1 Experiment Setups and Results
To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University (SYSU). It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5-inch PMT (HAMAMATSU CR284 with E2183-500 socket assembly). The crystal is wrapped with ESR (glued) and coupled with PMT using optical grease, finally coated with 3M tape (Fig. 6b). The readout system uses a Lecroy HDO4054A oscilloscope (12 bit, 10 GS/s, see Fig. 6c) for signal triggering, integration, and digitization. The benchmark detector is calibrated using ${}^{137}\text{Cs}$ ($7.81\times 10^{3}$ Bq) and ${}^{152}\text{Eu}$ ($3.55\times 10^{4}$ Bq) radioactive sources. Radioactive sources were positioned close to the detector, and data acquisition was performed for 1800 s. When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform over a 2 $\mu$s time window and records the charge value. The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig. 7. The spectra display the full-energy peaks of both isotopes. The energy resolution at 511 keV is calculated to be 11.5% by fitting the FWHM curve.
@@ III.2 Comparison and Discussion
Figure 8: The energy spectra of ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ in experiment (dots with error bars) and simulation (red line).
This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results. To test the reliability of the simulation results discussed in Section II, we conducted a detailed simulation of the benchmark CsI(Tl) detector using Geant4 with the same physics processes. The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup. Fig. 8 shows the spectra of calibrated experimental data and those recorded in simulation using ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ source. The energy resolution is 9.7% at 511 keV in the simulation, compared to 10.3% in the experiment (see Table 4), resulting in a relative difference of 6%. Note that the difference in the width and height of the full-energy peak between the MC and the data originates from the discrepancy in energy resolution. Other minor discrepancies in the spectra shape are mainly likely caused by the insufficient modeling of electronic noise and nonlinearity, natural backgrounds, materials around the detector, and the simplified optical processes in Geant4 [59]. This work focuses on a bench test of the CsI(Tl) crystal. However, important factors mentioned above that are difficult to simulate will be further studied in future prototype development efforts.
@@ IV Conclusions and Perspectives
However, the studies on accidental coincidence backgrounds from the beam and cosmic rays are currently idealized, and the effects of electronic noise and natural radioactive backgrounds have not yet been evaluated. A prototype of ECAL will be constructed for additional validation in progress in these areas. Further studies are anticipated to evaluate the performance enhancement of ECAL with large-area SiPM array readout. Upgrades of ECAL are expected with the integration of an inner tracker system. Moreover, future works will also focus on developing systems for cosmic-ray vetoing, electronic readout, trigger logic, and reconstruction algorithms.

### v4

@@ I Introduction
In light of significant advancements in muon beamlines and particle detector techniques nowadays, the Muonium-to-Antimuonium Conversion Experiment (MACE) is proposed to search for M-to-$\bar{\text{M}}$ conversion and has the potential to improve the upper limit by three orders of magnitude [18, 19]. With the construction of the China advanced NUclear physics research Facility (CNUF) and other local muon sources [20, 21, 22, 23], MACE is expected to be the first muon-CLFV experiment in China, offering unique perspectives compared to other related experiments, including those at colliders. The proposed MACE detector consists of a Michel electron Magnetic Spectrometer (MMS), a Positron Transport System (PTS), a Microchannel Plate (MCP) [24, 25], and an Electromagnetic Calorimeter (ECAL) to probe the conversion signal through final state coincidence. The specifications of muon beamline in MACE are expected to produce a surface muon flux on target of $10^{8}~{}\mu^{+}/$s, with a repetition rate of 30–50 kHz and a pulse width duty cycle below 20%. A schematic of the MACE apparatus is shown in Fig. 1.
@@
In the MACE concept, surface muons with a momentum of 26.3 MeV$/c$ are stopped at a production target where M is formed and then possibly diffuses into the vacuum and converts into $\bar{\text{M}}$. The decay products of $\bar{\text{M}}$ include a Michel electron with an energy up to 52.8 MeV and a positron with atomic-shell energy ($\langle E_{\text{k}}\rangle=13.5$ eV). The Michel electron will be reconstructed by MMS which is expected to achieve $10^{-4}$ level suppression of background in the ECAL end through preliminary simulation. On the other hand, the atomic-shell positron will be accelerated by an electric field and guided through the PTS. The transverse position of the $e^{+}$ is conserved during the transportation to MCP, with a flight time full width at half maximum (FWHM) of 6.9 ns. The cylindrical magnet with a field strength of 0.1 T outside the ECAL is used to constrain the positron beam spot after transportation. When the MCP, with a timing resolution of 500 ps, is hit by a positron, the ECAL will be triggered and measure the produced two back-to-back annihilation $\gamma$-rays. As a rare-process searching experiment, MACE requires high sensitivity for the annihilation $\gamma$-rays to discriminate the signal from backgrounds and retain as much event information as possible. Thus, an ECAL with good energy resolution and high efficiency is essential for the design of MACE.
@@ II.1 Overview
Calorimeters with inorganic scintillation crystals generally provide excellent energy resolution and thus have been widely used in particle physics experiments for decades. Commonly used scintillators in calorimeters include NaI(Tl), pure CsI, CsI(Tl), BGO, PbWO4, and others [43]. Table 1 summarizes the properties of these inorganic scintillators. The scintillation crystal selection for the MACE ECAL is based on the energy region of interest less than 1 MeV. NaI(Tl) was widely used in the past decades due to its low cost and high scintillation light output. Nonetheless, it is hygroscopic and has a relatively low density, making it less suitable for large-scale particle physics experiments at present. The problem of hygroscopy is identical for LaBr3(Ce). Moreover, considering the overall budget of this experiment, ultrafast and bright crystals such as LYSO(Ce), GAGG(Ce), or LaBr3(Ce) are not currently included but may be considered for potential future upgrades. The low light output of PbWO4 leads to an insufficient energy resolution in the low-energy regime. The same is true for pure CsI. In the MACS experiment, the energy resolution of the pure CsI calorimeter for single-$\gamma$ event and double-$\gamma$ events was 69% and 52%, respectively [44, 14]. Doping thallium (Tl) in pure CsI can significantly increase its light yield, making it one of the brightest scintillators known. The objective of MACE ECAL is to achieve excellent energy resolution for $\gamma$-rays produced in $e^{+}e^{-}$ annihilation. This requires optimizing the light yield of the scintillator, which is a key factor in precise energy measurements. Consequently, CsI(Tl) is chosen as the reference scintillator in this ECAL design.
@@ II.2 Optimization of Geometry Parameters
In the context of the MACE experiment, the signal positrons from $\bar{\text{M}}$ decay are transported through a solenoid into the region of the positron detection system. Some modules must be removed to ensure that the incoming signal positrons are not obstructed. Specifically, the beam windows are set with an upstream radius of 50 mm and a downstream radius of 5 mm. Consequently, 19 modules at the beam entrance and 1 module at the exit are removed, resulting in a total of 622 modules and achieving 97% coverage of the $4\pi$ solid angle. For an estimation of energy resolution, the class G4RadioactiveDecayPhysics is used to simulate the calibration of the Type-HEX06 module with radioactive sources, including 22Na, 137Cs, and 152Eu. Instead of the fully-covered SD in the previous simulation, the crystal is coupled with a 1.5-inch diameter PMT volume. An isotropic point source, wrapped in 1 mm thick plastic, is placed 5 cm away from the module. The energy dependence of FWHM is fit using Eq. 2:
@@
where the best-fit value of $a$, $b$ are $-7.4707$ and $2.7638$, respectively. The fit FWHM and energy resolution results of each full-energy peak are plotted in Fig. 4. For the Type-HEX06 modules, the energy resolution is 10.8% at 511 keV, which is 6–7 times better than the resolution achieved by the calorimeter used in the MACS experiment [14].
@@ II.3.1 M-to-$\bar{\text{M}}$ conversion signal
Figure 5: Spectra of a single Type-PEN module (blue line), a single Type-HEX module (green line) and reconstructed annihilation signals (red line) of M-to-$\bar{\text{M}}$ conversion by simulation.
@@ II.3.2 Beam-related backgrounds
The muon beam, produced by stopping $\pi^{+}$ on the surface of a target, often includes positrons or other particles with similar momentum [54]. These background particles can enter the ECAL chamber in the same manner as the signal positrons, potentially causing accidental coincidences between the MCP and the ECAL. Therefore, it is crucial to estimate their potential contribution.
@@ II.3.3 Cosmic-Ray Muon backgrounds
Cosmic-ray muons are another source of background in ECAL. The EcoMug generator [56] is integrated into the MACE offline software to simulate cosmic-ray muon events on various types of surfaces (plane, cylinder, and half-sphere) with accurate angle and momentum distributions. To evaluate the cosmic-ray muon background level in a potential future environment for the MACE detector, a hypothetical 18 m deep tunnel has been modeled in Geant4. A concrete volume, represented by G4Box with dimensions $50\times 50\times 18~{}\text{m}^{3}$, is created, with an air-filled cylindrical region, represented by G4Tubs, subtracted from its bottom to form a tunnel with a 1 m radius and a 10 m length. ECAL is positioned at the center of the tunnel, surrounded by a copper solenoid and an iron shield. Cosmic-ray muons are emitted from a $50\times 50~{}\text{m}^{2}$ plane source at sea level. With a rate of 129 Hz, cosmic-ray muon events equivalent to 1 hour of exposure are generated.
The simulation results indicate that cosmic-ray muons still contribute to the background in the ECAL even after passing through the concrete and shielding. This includes some events where $e^{+}e^{-}$ annihilation produces 511 keV $\gamma$ rays. After applying coincidence selection, no events remain, resulting in a background rate of less than 2/2/2 /yr when considering coincidences with the Magnetic Spectrometer. Typically, a veto system can be installed outside ECAL to further reject cosmic-ray muon background [6, 57].
II.3.4 Radioactive contamination background
It is worth mentioning that studies have reported the radioactive contamination of scintillation materials (e.g., 137Cs in CsI), which may greatly affect low-counting experiments. In the context of the MACE PDS, if the MCP is triggered by its dark counts while the radionuclides within the ECAL materials decay simultaneously, a false coincidence may occur. Based on the provided data, a single CsI(Tl) module in the ECAL will have a 662 keV $\gamma$-ray background counting rate up to 0.07 Hz [58]. A preliminary estimation of the accidental coincidence rate can be conducted by
$\begin{gathered}n=2\tau n_{\text{MCP}}n_{\text{ECAL}}~{},\\ n_{\text{ECAL}}=2\tau n_{\text{mod1}}n_{\text{mod2}}\times N_{\text{mod}}/2~{},\end{gathered}$ (3)
where $\tau$ is the coincidence resolving time which is set to be CsI(Tl) decay time, $n_{\text{MCP}}$ is the dark count rate of the MCP (commonly $\sim 1$ kHz), $n_{\text{ECAL}}$ is the accidental coincidence of two modules caused by the natural background detected by the ECAL and $N_{\text{mod}}$ is the total module numbers. Eventually, the rate is calculated to be 0.19/0.19/0.19 /yr. This background, however, falls outside the $3\sigma$ interval of the signal of interest. If the event selection criteria are applied, the background caused by dark count and natural radioactivity can be neglected.
@@ III.1 Experiment Setups and Results
To characterize the parameters of the CsI(Tl) scintillator, a benchmark detector was constructed and calibrated using various radioactive sources at Sun Yat-sen University (SYSU). It consists of a $3\times 3\times 8~{}\text{cm}^{3}$ CsI(Tl) crystal and a 1.5-inch PMT (HAMAMATSU CR284 with E2183-500 socket assembly). The crystal is wrapped with ESR (glued) and coupled with PMT using optical grease, finally coated with 3M tape (Fig. 6b). The readout system uses a Lecroy HDO4054A oscilloscope (12 bit, 10 GS/s, see Fig. 6c) for signal triggering, integration, and digitization. The benchmark detector is calibrated using ${}^{137}\text{Cs}$ ($7.81\times 10^{3}$ Bq) and ${}^{152}\text{Eu}$ ($3.55\times 10^{4}$ Bq) radioactive sources. Radioactive sources were positioned 5 cm away from the detector, and data acquisition was performed for 1800 s. When the PMT pulse signal exceeds the threshold, the oscilloscope triggers, integrates the waveform over a 2 $\mu$s time window, and records the charge value. The calibration used the 662 keV peak from ${}^{137}\text{Cs}$ and the 40 keV, 122 keV, 244 keV, 344 keV, and 1408 keV peaks from ${}^{152}\text{Eu}$, as shown in Fig. 7. The spectra display the full-energy peaks of both isotopes. The energy resolution at 511 keV is calculated to be 11.5% by fitting the FWHM curve.
@@ III.2 Comparison and Discussion
Figure 8: The energy spectra of ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ in the experiment (dots with error bars) and the simulation (red line).
This study primarily relies on Geant4 simulation for the design of the MACE calorimeter, requiring verification of the reliability of various physical processes in the simulation and a comparison with experimental results. To test the reliability of the simulation results discussed in Section II, we conducted a detailed simulation of the benchmark CsI(Tl) detector using Geant4 with the same physics processes. The CsI(Tl) crystal, Teflon reflector, optical grease, and CR284 PMT volume are modeled identically to the experimental setup. The simulation takes into account the natural background contributions from 238U and 40K. Fig. 8 shows the spectra of calibrated experimental data and those recorded in simulation using ${}^{22}\text{Na}$ and ${}^{137}\text{Cs}$ source. The energy resolution is 9.7% at 511 keV in the simulation, compared to 10.3% in the experiment (see Table 4), resulting in a relative difference of 6%. Note that the difference in the width and height of the full-energy peak between the MC and the data originates from the discrepancy in energy resolution. Other minor discrepancies in the spectra shape are mainly likely caused by the insufficient modeling of electronic noise and nonlinearity, natural backgrounds, materials around the detector, and the simplified optical processes in Geant4 [60]. This work focuses on a bench test of the CsI(Tl) crystal. Therefore, important factors mentioned above that are difficult to simulate will be further studied in future prototype development efforts.
@@ IV Conclusions and Perspectives
A prototype of ECAL will be constructed for additional validation in progress in these areas. Further studies are anticipated to evaluate the performance enhancement of ECAL with large-area SiPM array readout. Upgrades of ECAL are expected with the integration of an inner tracker system. Moreover, future works will also focus on developing systems for cosmic-ray vetoing, electronic readout, trigger logic, and reconstruction algorithms.